OpenStudy (anonymous):

Solve 2x + 5y = 10 for y

OpenStudy (anonymous):

y = 2 - 2x/5

OpenStudy (anonymous):

y = 2x + 2 y = –2x + 10 y = –2 over 5x + 2 y = 2 over 5x + 10 which one of these though?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Isolate y, y = (10-2x)/5 = 2 - 2/5 x

OpenStudy (anonymous):

Should be the 3rd one

OpenStudy (anonymous):

wow, thanks :)